OpenStudy (anonymous):

how would you solve 8 raised to the -2/3 power?

OpenStudy (anonymous):

\[8=2^{3}\] use this...

OpenStudy (anonymous):

the -2 means it's 1 over 8 squared.. or (1/8)^2 then take the cube root

OpenStudy (mathteacher1729):

In three parts: 1) 8 to the 1/3rd power (or the 3rd root of 8) 2) The answer from step (1) squared 3) 1 over the answer from step (2)

OpenStudy (anonymous):

\[(2^{3})^{-2/3}=2^{-2/3 *3}=2^{-2}=1/4\]

OpenStudy (anonymous):

Inik, I think I pretty much have it, but do you just cancel out both of those 3's or what?

OpenStudy (anonymous):

yes 3*(-2/3)=-2